OpenStudy (anonymous):

3. Graph –3x^2 + 12y^2 = 84. What are the domain and range?

OpenStudy (anonymous):

its an hyperbola. Domain : \[R^{2}\] Range: \[-\sqrt{7} \le y \le \sqrt{7}\]